# Occupied Processes: Going with the Flow

### v4

@@ 5.2.1 The LOV Model
      Update $O^{\kappa}$$\quad O^{\kappa}_{m}(\omega_{j})\leftarrow O^{\kappa}_{m}(\omega_{j})+e^{\kappa t_{n}}\delta t$, $X_{t_{n}}(\omega_{j})\in C_{m}$
@@
      $\hat{O}^{\kappa}_{m}(\omega_{j})=\frac{\sum_{j^{\prime}}\psi_{j,j^{\prime}}O^{\kappa}_{m}(\omega_{j^{\prime}})}{\sum_{j^{\prime}}\psi_{j,j^{\prime}}},\quad\psi_{j,j^{\prime}}=\psi(X_{t_{n}}(\omega_{j})-X_{t_{n}}(\omega_{j^{\prime}}))$
@@
Figure 7: Simulated price path and volatility in the Local Occupied Volatility (LOV) model.
Figure 8: Local occupied volatility compared with local volatility. The red curve is the local sensitivity function $x\mapsto\ell(x/X_{t})$.
5.3 Neural Calibration and Occupation Sensitivities
To gain insights into the LOV model, let us go back to the general occupied volatility (64) and parameterize $\sigma$ by a feedforward neural network, that is
$\sigma(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})\approx\sigma^{\theta}(O_{t}^{\kappa},X_{t})\mathrel{\mathop{\ordinarycolon}}=\Phi(O_{t}^{\kappa},X_{t};\theta),\quad\Phi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}^{M}\times\operatorname{\mathbb{R}}_{+}\times\Theta\to\operatorname{\mathbb{R}}_{+},$ (76)
with the occupation times $O_{t}^{\kappa}=(\operatorname{\mathcal{O}}_{t}^{\kappa}(C_{1}),\ldots,\operatorname{\mathcal{O}}_{t}^{\kappa}(C_{M}))\in\operatorname{\mathbb{R}}^{M}$. The asset price process solves therefore the following neural OSDE,
$\frac{dX_{t}^{\theta}}{X_{t}^{\theta}}=(r-q)dt+\sigma^{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})dW_{t},\qquad dO_{t,m}^{\kappa,\theta}=\mathds{1}_{C_{m}}(X_{t}^{\theta})e^{\kappa t}dt.$ (77)
The parameters $\theta\in\Theta$ are then calibrated so as to match market quotes of vanilla call and put options; see Algorithm 2. Note that a single neural network is used in (76) for all time points, which presents present significant benefits in terms of computational cost and smoothness in the time variable; see [73]. The optimization is also done on the full horizon at all once, often called global-in-time approach [60].
Neural networks offer an effective and flexible way to parametrize high-dimensional functions, at the cost of losing a clear interpretation of the parameters. Nevertheless, one way to understand the link between input (spot and occupation times) and output (volatility) is to compute sensitivities of the neural network, readily available through automatic differentiation; see e.g., [36, Chapter 5]. In our context, we can extract the partial derivatives of variance with respect to the occupation times $\operatorname{\mathcal{O}}_{t}^{\kappa,\theta}(C_{m})$, and retrieve the local sensitivity function $\ell$ in the LOV model (74)-(75) through
$\gamma_{t}^{\kappa}\ell(t,X_{t}^{\theta},x)=\delta_{\mathfrak{o}}\sigma^{2}(\operatorname{\mathcal{O}}_{t}^{\kappa,\theta},X_{t})(x)\approx\partial_{m}(\sigma^{\theta})^{2}(O_{t}^{\kappa,\theta},X_{t}^{\theta}),\quad x\in C_{m},$ (78)
where $\partial_{m}$ denotes the partial derivative with respect to $O_{m}$. Illustrations of $x\mapsto\ell(t,X_{t},x)$ will be given in Section 5.3.3.
@@ 5.3.1 Calibration
We then train the parameters of the neural volatility $\sigma^{\theta}$ in (77) so that the model prices,
@@
As the Vega becomes arbitrarily small for out-of-the-money options and short maturities, we also cap the weights by setting instead $(\vartheta^{\text{\tiny mid}}_{i}\vee\varepsilon)^{-1}$ in (79) with, e.g., $\varepsilon=10^{-2}$; see also [5]. Moreover, the bid-ask ratio $\frac{\sigma^{\text{\tiny bid}}_{i}}{\sigma^{\text{\tiny ask}}_{i}}$ gives greater weight to liquid contracts, i.e., those with tighter percentage bid-ask spread $1-\frac{\sigma^{\text{\tiny bid}}_{i}}{\sigma^{\text{\tiny ask}}_{i}}$. Although one could directly use the implied volatility errors $|\sigma^{\theta}_{i}-\sigma^{\text{\tiny mid}}_{i}|$ in the loss function, this would necessitate the computation of the model implied volatilities $(\sigma^{\theta}_{i})_{i=1}^{I}$ at every training step, which becomes costly for large $I$. Note in contrast that the bid-ask ratios $\frac{\sigma^{\text{\tiny bid}}_{i}}{\sigma^{\text{\tiny ask}}_{i}}$ can be computed offline.
@@
    $\operatorname{\mathcal{L}}(\theta)=\Big{(}\frac{1}{I}\sum_{i=1}^{I}|w_{i}(\pi^{\theta}_{i}-\pi^{\text{\tiny mid}}_{i})|^{2}\Big{)}^{1/2},\quad w_{i}=(\vartheta^{\text{\tiny mid}}_{i})^{-1}\cdot\frac{\sigma^{\text{\tiny bid}}_{i}}{\sigma^{\text{\tiny ask}}_{i}},$ (79)
@@ 5.3.2 Market Data and Results
$2025/\text{mm}/\text{dd},\quad\text{mm}/\text{dd}\in\{03/21,04/18,05/16,06/20,07/18,08/15\}.$
@@
We sample $2^{8}$ pairs of antithetic paths (hence $J=2^{9}$) for each training iteration, which is increased to $2^{12}$ pairs after $1,000$ epochs. In Step I., the occupied SDE is discretized using $\delta t=\frac{1}{252}$ (daily time step) and $M=2^{6}-1=63$ bins centered at equally spaced nodes between $X_{0}^{T}\pm 2\sigma_{\text{\tiny ATMF}}^{T}$, where $X_{0}^{T}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[X_{T}]\approx 6,000$ is the market implied forward price from the six-month options (hence $T\approx 0.5$) and $\sigma_{\text{\tiny ATMF}}^{T}$ is the associated at-the-money forward implied volatility. We also choose an exponential time parameter of $\kappa=12$ (one month horizon). The neural network $\Phi$ contains two hidden layers containing $2^{6}=64$ nodes, with the ReLU and softplus activation function for the hidden and output layers, respectively.
@@
5.3.3 Path Dependence and Occupation Sensitivities
@@
Fig. 11 Displays the volatility values $\sigma^{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})$ against the spot $X_{t}^{\theta}$ for $2^{10}$ trajectories on the full calibration horizon ($t\approx 0.5$). Observe that the dots do not lie on the graph of $\hat{\sigma}^{\theta}(t,\cdot)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma^{2}(\operatorname{\mathcal{O}}_{t}^{\kappa,\theta},X^{\theta}_{t})|X_{t}^{\theta}=\cdot]^{1/2}$, confirming the presence of path dependence in $\sigma^{\theta}$. Moreover, the volatility level typically increases as $X_{t}^{\theta}$, in line with the leverage effect. Note also that $\hat{\sigma}^{\theta}$ is a smooth approximation of the local volatility $\sigma_{\text{loc}}$ since the neural occupation volatility has been calibrated to vanilla options.
Figure 11: Neural volatility $\sigma^{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})$ versus spot $X_{t}^{\theta}$ for $2^{10}$ trajectories.
Back to the LOV model, we can finally visualize the local sensitivity function $\ell(t,X_{t},x)$ computed through formula (78). Fig. 12 displays three scenarios where the spot as of 2025/08/15 ends respectively below, close to, and above the initial value $X_{0}$. In every scenario, the sensitivity (in red) is positive for smaller $x$ and flips sign around the spot $X_{t}$, in line with expectation. Indeed, recall that $\ell(t,X_{t},x)$ describes the change in variance when the asset jumps to level $x$ for a short time period. If $x$ is less than the spot, we then expect the variance (or equivalently volatility) to increase, i.e., $\ell>0$, and vice versa as confirmed in Fig. 12.
Figure 12: Occupation sensitivities (red) and projected measure $\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\operatorname{\mathcal{O}}_{t}^{\kappa}|X_{t}]$ (blue).
@@ 6 Forward Occupation Models
This section introduces a generalization of forward variance models [10, 49] that uses the information content of vanilla options not only through the forward variance curve, but through all forward occupation times. [^5] I would like to thank an anonymous referee for suggesting this promising direction. Consider a standard Brownian motion $W=(W^{1},\ldots,W^{d})$ in $\operatorname{\mathbb{R}}^{d}$, $d\geq 1$, and denotes its natural filtration by $(\operatorname{\mathcal{F}}_{t})$. Assume that the log price $X=\log S$ of an asset has dynamics
@@
where $\tau\geq 0$ depends on $t,T$. In Bergomi models [10], $\tau=T$ with $\eta(t,T,\xi_{t}^{T})$ linear in $\xi_{t}^{T}$, while in affine forward variance models [49], $\tau=t$, namely $\eta(t,T,\xi_{t}^{\tau})=\eta(t,T,V_{t})$. See Examples 15 and 16. Given a variance swap that delivers the realized variance on $[0,T]$, then the value of its (unannualized) floating leg at time $t\in[0,T]$ can then be written as
@@
$\xi_{t}^{T}\hookrightarrow\mathfrak{o}_{t}^{T}(A)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{t}[\mathds{1}_{A}(X_{T})V_{T}],\quad A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}),$ (82)
@@ 6.1 Dynamics of Forward Occupation Flows
$\mathfrak{o}_{t}^{T}(A)=\mathfrak{o}_{0}^{T}(A)+\int_{0}^{t}\eta^{T}_{u}(A)\cdot dW_{u},\quad t\leq T,$ (84)
@@
$df(\operatorname{\mathcal{O}}^{T}_{t})$ $=(\delta_{\mathfrak{o}}f(\operatorname{\mathcal{O}}_{t}^{T})\cdot H_{t}^{T})\cdot dW_{t}+\frac{1}{2}\delta_{\mathfrak{o}\mathfrak{o}}f(\operatorname{\mathcal{O}}_{t}^{T})\cdot(H_{t}^{T})^{\otimes 2}dt,\quad f\in\operatorname{\mathcal{C}}^{2}(\operatorname{\mathcal{M}}),$ (86)
@@
$\varphi(\operatorname{\mathcal{O}}_{T})=(\operatorname{\mathcal{O}}_{T}(A^{\pm})-K)^{+},\quad A^{-}=(-\infty,X_{0}],\quad A^{+}=(X_{0},\infty).$
@@
$\frac{1}{\Delta}w\cdot(\operatorname{\mathcal{O}}_{T}^{T+\Delta}-\operatorname{\mathcal{O}}_{T})=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{T}\Big{[}\frac{1}{\Delta}\int_{T}^{T+\Delta}w(X_{s})V_{s}ds\Big{]},\qquad w\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}},\quad\Delta>0.$
@@
$H_{t}^{T}=\delta_{X_{t}}\eta_{0}\sqrt{V_{t}}K(t,T),\quad K(t,T)=\int_{t}^{T}\kappa(s-t)ds\in\operatorname{\mathbb{R}}^{d}.$ (88)

### v5

@@ 5.2.2 Simulation
      Update $O^{\kappa}$$\qquad O^{\kappa}_{m}(\omega_{j})\leftarrow O^{\kappa}_{m}(\omega_{j})+e^{\kappa t_{n}}\delta t$, $X_{t_{n}}(\omega_{j})\in C_{m}$
@@
      $\hat{O}^{\kappa}_{m}(\omega_{j})=\frac{\sum_{j^{\prime}}\psi_{j,j^{\prime}}O^{\kappa}_{m}(\omega_{j^{\prime}})}{\sum_{j^{\prime}}\psi_{j,j^{\prime}}},\qquad\psi_{j,j^{\prime}}=\psi(X_{t_{n}}(\omega_{j})-X_{t_{n}}(\omega_{j^{\prime}}))$
@@
5.3 Market Implied Occupation Sensitivity
[…]
$\sigma(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})\approx\sigma_{\theta}(O_{t}^{\kappa},X_{t})\mathrel{\mathop{\ordinarycolon}}=\Phi(O_{t}^{\kappa},X_{t};\theta),\qquad\Phi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}^{M}\times\operatorname{\mathbb{R}}_{+}\times\Theta\to\operatorname{\mathbb{R}}_{+},$ (77)
with the occupation times $O_{t}^{\kappa}=(\operatorname{\mathcal{O}}_{t}^{\kappa}(C_{1}),\ldots,\operatorname{\mathcal{O}}_{t}^{\kappa}(C_{M}))\in\operatorname{\mathbb{R}}^{M}$. The asset price process thus evolves according to the following neural occupied SDE,
$\frac{dX_{t}^{\theta}}{X_{t}^{\theta}}=(r-q)dt+\sigma_{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})dW_{t},\hskip 17.00024ptdO_{t,m}^{\kappa,\theta}=\mathds{1}_{C_{m}}(X_{t}^{\theta})e^{\kappa t}dt.$ (78)
The parameters $\theta\in\Theta$ are then calibrated to market quotes of vanilla call and put options (see Algorithm 2). Note that a single neural network is used in (77) for all time points, which presents present significant benefits in terms of computational cost and smoothness in the time variable; see [74]. The optimization is also done on the full horizon at all once, often called global-in-time approach [60].
Neural networks offer an effective and flexible way to parametrize high-dimensional functions, at the cost of losing a clear interpretation of the parameters. Nevertheless, one way to understand the link between input (spot and occupation times) and output (volatility) is to compute sensitivities of the neural network, readily available through automatic differentiation; see e.g., [36, Chapter 5]. In our context, we can extract the partial derivatives of variance with respect to the occupation times $\operatorname{\mathcal{O}}_{t}^{\kappa,\theta}(C_{m})$, and retrieve the implied sensitivity $\ell^{\operatorname{\mathbb{Q}}}$ through
$\gamma_{t}^{\kappa}\ell^{\operatorname{\mathbb{Q}}}(t,X_{t}^{\theta},x)=\delta_{\mathfrak{o}}\sigma^{2}(\operatorname{\mathcal{O}}_{t}^{\kappa,\theta},X_{t})(x)\approx\partial_{m}\sigma_{\theta}^{2}(O_{t}^{\kappa,\theta},X_{t}^{\theta}),\qquad x\in C_{m},$ (79)
where $\partial_{m}$ denotes the partial derivative with respect to $O_{m}$. Illustrations of $\ell^{\operatorname{\mathbb{Q}}}$ will be given in Section 5.3.3.
@@ 5.3.1 Calibration
We then train the parameters of the neural volatility $\sigma_{\theta}$ in (78) so that the model prices,
@@
As the Vega becomes arbitrarily small for out-of-the-money options and short maturities, we also cap the weights by setting instead $(\vartheta^{\text{\tiny mid}}_{i}\vee\varepsilon)^{-1}$ in (80) with, e.g., $\varepsilon=10^{-2}$; see also [5]. Moreover, the bid-ask ratio $\frac{\sigma^{\text{\tiny bid}}_{i}}{\sigma^{\text{\tiny ask}}_{i}}$ gives greater weight to liquid contracts, i.e., those with tighter percentage bid-ask spread $1-\frac{\sigma^{\text{\tiny bid}}_{i}}{\sigma^{\text{\tiny ask}}_{i}}$. Although one could directly use the implied volatility errors $|\sigma^{\theta}_{i}-\sigma^{\text{\tiny mid}}_{i}|$ in the loss function, this would necessitate the computation of the model implied volatilities $(\sigma^{\theta}_{i})_{i=1}^{I}$ at every training step, which is prohibitively costly for large $I$. Note in contrast that the bid-ask ratios $\frac{\sigma^{\text{\tiny bid}}_{i}}{\sigma^{\text{\tiny ask}}_{i}}$ can be computed offline.
@@
    $\operatorname{\mathcal{L}}(\theta)=\Big(\frac{1}{I}\sum_{i=1}^{I}|w_{i}(\pi^{\theta}_{i}-\pi^{\text{\tiny mid}}_{i})|^{2}\Big)^{1/2},\qquad w_{i}=(\vartheta^{\text{\tiny mid}}_{i})^{-1}\cdot\frac{\sigma^{\text{\tiny bid}}_{i}}{\sigma^{\text{\tiny ask}}_{i}},$ (80)
@@ 5.3.2 Market Data and Results
$2025/\text{mm}/\text{dd},\qquad\text{mm}/\text{dd}\in\{03/21,04/18,05/16,06/20,07/18,08/15\}.$
@@
We sample $2^{8}$ pairs of antithetic paths (hence $J=2^{9}$) for each training iteration, which is increased to $2^{12}$ pairs after $1,000$ epochs. In Step I., the occupied SDE is discretized using $\delta t=\frac{1}{252}$ (daily time step) and $M=2^{6}-1=63$ bins centered at equally spaced nodes between $X_{0}^{T}\pm 2\sigma_{\text{\tiny ATMF}}^{T}$, where $X_{0}^{T}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[X_{T}]\approx 6,000$ is the market implied forward price from the six-month options (hence $T\approx 0.5$) and $\sigma_{\text{\tiny ATMF}}^{T}$ is the associated at-the-money forward implied volatility. We also choose calendar time parameter ($\kappa=0$) so the calibrated model could be used to price/manage exotic options according to the unified framework Fig. 3. The neural network $\Phi$ contains two hidden layers containing $2^{6}=64$ nodes, with the ReLU and softplus activation function for the hidden and output layers, respectively.
@@
5.3.3 Implied Sensitivity Function
@@
Fig. 11 Displays the volatility values $\sigma_{\theta}(O_{t}^{\kappa,\theta},X_{t}^{\theta})$ against the spot $X_{t}^{\theta}$ for $2^{10}$ trajectories on the full calibration horizon ($t\approx 0.5$). Observe that the dots do not lie on a one-dimensional manifold, confirming the presence of path dependence in $\sigma_{\theta}$. Moreover, the volatility level typically increases as $X_{t}^{\theta}$, in line with the leverage effect. Note also that $\hat{\sigma}^{\theta}$ is a smooth approximation of the local volatility $\sigma_{\text{loc}}$ since the neural occupation volatility has been calibrated to vanilla options.
Figure 11: Neural and local volatility versus spot. $2^{10}$ samples.
[…]
Figure 12: Implied sensitivity function $\ell^{\operatorname{\mathbb{Q}}}(t,X_{t},\cdot)$ (yellow) and projected occupation measure $\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\operatorname{\mathcal{O}}_{t}^{\kappa}|X_{t}]$ (purple).
5.4 Joint Calibration
To better pinpoint the sensitivity function, further work includes the calibration of occupied volatility to other instruments, such as VIX, exotic, or American options. Indeed, while the local volatility function guarantees perfect calibration to European vanilla options, the sensitivity function $\ell^{\operatorname{\mathbb{Q}}}$ can be tuned so as to match the price of other instruments, as summarized below:
$\sigma^{2}(\mathcal{O}_{t}^{\kappa},X_{t})=\underbrace{\sigma_{\text{loc}}^{2}(t,X_{t})}_{\text{European vanillas}}+\ \gamma_{t}^{\kappa}\int_{\operatorname{\mathbb{R}}_{+}}\ \underbrace{\ell^{\operatorname{\mathbb{Q}}}(t,X_{t},x)}_{\begin{subarray}{c}\text{VIX, exotic}\\ \text{American}\end{subarray}}\ (\mathcal{O}_{t}^{\kappa}-\hat{\mathcal{O}}^{\kappa}_{t})(dx).$
Also, from our experiments, the sensitivity function seems critical to build a model consistent with American-style vanilla options on U.S. single names, where one should typically go beyond local volatility for this task. One interesting case arises when the underlying does not pay dividends, such as Amazon.com, Inc. (AMZN). Indeed, the price of American call options coincides with their European counterparts [70], so the local volatility is uniquely characterized by the call options through Dupire’s formula. On the other hand, assuming positive interest rates, American put options carry an early exercise premium, and the local volatility model is not guaranteed to perfectly replicate the market prices of these options. The author is currently investigating the application of the LOV model to this joint American/European calibration problem.
@@ 6 Forward Occupation Models
This section introduces a generalization of forward variance models [10, 49] that uses the information content of vanilla options not only through the forward variance curve, but through all forward occupation times. [^5] I would like to thank an anonymous referee for suggesting this promising direction. Consider a standard Brownian motion $W=(W^{1},\ldots,W^{d})$ in $\operatorname{\mathbb{R}}^{d}$, $d\geq 1$, and denote its natural filtration by $(\operatorname{\mathcal{F}}_{t})$. Assume that the log price $X=\log S$ of an asset has dynamics
@@
where $\tau\geq 0$ depends on $t,T$. In Bergomi models [10], $\tau=T$ with $\eta(t,T,\xi_{t}^{T})$ linear in $\xi_{t}^{T}$, while in affine forward variance models [49], $\tau=t$, namely $\eta(t,T,\xi_{t}^{\tau})=\eta(t,T,V_{t})$. See Examples 17 and 18. Given a variance swap that delivers the realized variance on $[0,T]$, then the value of its (unannualized) floating leg at time $t\in[0,T]$ can then be written as
@@
$\xi_{t}^{T}\hookrightarrow\mathfrak{o}_{t}^{T}(A)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{t}[\mathds{1}_{A}(X_{T})V_{T}],\qquad A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}),$ (83)
@@ 6.1 Dynamics of Forward Occupation Flows
$\mathfrak{o}_{t}^{T}(A)=\mathfrak{o}_{0}^{T}(A)+\int_{0}^{t}\eta^{T}_{u}(A)\cdot dW_{u},\qquad t\leq T,$ (85)
@@
$df(\operatorname{\mathcal{O}}^{T}_{t})$ $=(\delta_{\mathfrak{o}}f(\operatorname{\mathcal{O}}_{t}^{T})\cdot H_{t}^{T})\cdot dW_{t}+\frac{1}{2}\delta_{\mathfrak{o}\mathfrak{o}}f(\operatorname{\mathcal{O}}_{t}^{T})\cdot(H_{t}^{T})^{\otimes 2}dt,\qquad f\in\operatorname{\mathcal{C}}^{2}(\operatorname{\mathcal{M}}),$ (87)
@@
$\varphi(\operatorname{\mathcal{O}}_{T})=(\operatorname{\mathcal{O}}_{T}(A^{\pm})-K)^{+},\qquad A^{-}=(-\infty,X_{0}],\qquad A^{+}=(X_{0},\infty).$
@@
$\frac{1}{\Delta}w\cdot(\operatorname{\mathcal{O}}_{T}^{T+\Delta}-\operatorname{\mathcal{O}}_{T})=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{T}\Big[\frac{1}{\Delta}\int_{T}^{T+\Delta}w(X_{s})V_{s}ds\Big],\hskip 17.00024ptw\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}},\qquad\Delta>0.$
@@
$H_{t}^{T}=\delta_{X_{t}}\eta_{0}\sqrt{V_{t}}K(t,T),\qquad K(t,T)=\int_{t}^{T}\kappa(s-t)ds\in\operatorname{\mathbb{R}}^{d}.$ (89)